# Trapping polar molecules by surface acoustic waves

### v1

@@ IV.2 Cooperative shielding
We first consider a simplified case that $\gamma_{nj}\equiv\gamma$, and then consider a general case. As shown in Fig. 4(b), where $c_{n}(T)$ is for $\gamma\neq 0$ and $c_{n}^{\prime}(T)$ is for $\gamma=0$ [67], then the effect of $\gamma$ can be evaluated with $\max_{n}\left|\left|c_{n}(T)\right|^{2}-\left|c_{n}^{\prime}(T)\right|^{2}\right|$. The comparisons in Figs. 4(c) and (d) reveal that the long range interactions among trapped polar molecules can be cancelled out when the number of trapped sites is large based on the designed initial condition.
@@
where $c_{n}(t)$, as a generalization of $p_{n}$ in Eq. (15), represents the amplitude that the $n$th site is occupied. For the single-layer trapped polar molecules with a single SAW layer and externa electrical fields, $N$ can be finite.
When the polar molecules are trapped with two SAW layers as in Fig. 2(a), $N$ can be infinitely large. Given the Hamiltonian as Eq. (23), the evolution of the amplitudes can be written according to the Schrödinger equation $|\dot{\psi}(t)\rangle=-iH_{N}|\psi(t)\rangle$ as [67]
@@
with $\hbar=1$. Then in Fig. 5, we compare the dynamics of Eq. (25) with different number of sites and different initial conditions. The simulations clarify that long-range interactions can affect the dynamics of the lattices, but it is related to the initial condition and number of the sites. If the lattices are initially random or uniformly occupied, the influence of long-range interactions is much smaller or better shielded than the case that a small subset of lattices are initially occupied.
Figure 5: $||c_{n}(T)|^{2}-|c_{n}^{\prime}(T)|^{2}|$ varies with the number of cites $N$, where $c_{n}(T)$ represents the terminal values in Eq. (25) with $\gamma_{nj}=0$ at $T=10$s, and $c_{n}^{\prime}(T)$ represents that when $\gamma_{nj}\neq 0$. In all the three simulations, $V=1$ and $U_{0}=0.1$. In (a-c), $\gamma_{nj}\equiv\gamma=2$, and in (d-f), $\gamma_{nj}=V|n-j|^{-3}$. In (a) and (d), $c_{n}(0)=1/\sqrt{N}$. In (b) and (e), $c_{n}(t)$ and $c_{n}^{\prime}(t)$ are randomly initialized. In (c) and (f), $c_{1}(0)=c_{2}(0)=1/\sqrt{2}$ and $c_{n}(0)=0$ for $n>2$.
@@ V Bose-Hubbard model for trapped molecules
$\approx\Delta\left(\frac{\Delta^{2}-1}{2m_{0}}+\tilde{B}_{0}e^{-kz}+\delta_{J}\right)e^{-\Delta^{2}}$
@@
where $\delta_{\epsilon}$ is a bounded perturbation similarly as in Eq. (30) within a short time scale. In the following, we first study a simplified case with $\delta_{J}=\delta_{\epsilon}=0$ and then generalize to the case that $\delta_{J}=\delta_{\epsilon}\neq 0$.
@@ V.1 Influence by $n_{0}$, uncertainties and thermal effects
Figure 7: (a) Phase transition boundaries when $n_{0}=1,4,7,10,13,16$ and $f_{n_{0}}=0.085786,0.027864,\cdots,0.0075775$, respectively. (b) The parameters of the Bose-Hubbard model affected by $\delta_{J}\in[-5,5]$ and $\delta_{\epsilon}\in[-100,100]$. (c) The phase transition between superfluid (red) and Mott insulator (blue) evaluated by $J/U$ with the uncertain parameters in (b). (d) $J/U$ affected by the wave number $k$ when $N=5$ and $\delta_{J}\in[-5,5]$. In all the simulations, $\tilde{B}_{0}=100$, the distance between two nearest neighbor trapped polar molecules is $0.005/N$m, and $k=50$ in (a)-(c).
@@ V.2 Influence by the wave number
Figure 9: The influence on the phase transition by the wave number $k$ of the acoustic wave. (a)-(c) represent the phase transition between the superfluid (red region) and Mott insulator (blue region) when $n_{0}=1$ for different $k$. (d)-(f) represent the phase transition boundaries with $n_{0}=1,4,7,10,13,16$, and different colors represent the regions for the superfluid state separated by the boundaries with different $n_{0}$ compared with the white insulator region. In all the simulations, $B_{0}=100$, the distance between two nearest neighbor trapped polar molecules is $0.005/N$m.
@@ VI Conclusion
We have proposed a method to trap polar molecules with the electrical field induced by surface acoustic waves. Assisted by external electrical fields, the polar molecules can be trapped and arranged into single or multi-parallel layers away from the piezoelectric material. Depending on the uniformity of the layers in the longitudinal direction, the attractive interactions among the molecules in each layer can be different, which can further affect the binding energy of the trapped polar molecule array. For a single layer of trapped polar molecules, the final steady distribution of the molecules can be affected by the trapping approach with finite or infinite lattice sites. The trapped polar molecules can be used to construct the lattice based Bose-Hubbard model. The phase transition between the superfluid and Mott insulator can be controlled by designing the values and spatial distributions of the external electrical force. The advantages of SAW for trapping polar molecules are that its induced electrical field in the open space naturally satisfies the condition of stability for trapping a polar particle, the field can be real-time modulated by the electrical field applied to the input IDT, and arbitrary three-dimensional polar molecule lattices can in principle be trapped by tuning another applied electrical force. Our proposal can be more efficient compared with traditional trapping methods with electrical force such as in Refs. [21, 22, 23, 24]. Besides, the exponential decrease of the electrical field produced by SAW along the vertical direction in the open space can induce the transition between Mott insulator and superfluid, and the overall architecture can be fabricated on-chip with extensive possible expansions. From the perspective of experimental realization, the deceleration of the molecule gas has been realized in Refs. [20, 53, 90], the on-chip design of the surface acoustic wave resonator has been demonstrated in Ref. [54], and the method of shaping electrical fields in the open space between two SAW layers has been introduced in Refs. [21, 55]. These technologies support the viability of the proposed trapping method. Our proposal has potential applications in the construction of superfluid and Mott insulator [91, 92], phase transitions [93, 94, 95], lattice QED [96, 75], as well as quantum information processing [97]. This might open an easy way to hybridize molecules with solid state quantum devices, in contrast to the trapping molecules using optical lattices.
Acknowledgement.— Y.X.L. is supported by the NSFC under Grants No. 12374483 and No. 92365209. R.B.W. is supported by the NSFC under Grants No. 61833010 and No. 62173201.
@@ A.1 Solutions in isotropic materials
Then the vector format of the displacement $\mathbf{u}=\left[u_{1},u_{2},u_{3}\right]^{T}$ can be represented as [98]
@@
where $\bar{U}$, $\bar{V}$ and $\bar{W}$ are constants determined by $q$ representing the possible decay in the direction of $x_{3}$, $k$ is the wave number, $v$ is the velocity of the acoustic wave, $l=\cos\theta$, $m=\sin\theta$, $\theta$ is the angle between the propagating direction of the wave front with the $x_{3}$-axis. Thus $l$ and $m$ can represent the propagating direction of the acoustic wave. Take Eq. (45) into Eqs. (42,43,44), then we have $\mathbf{P}(q)\left[\bar{U}~{}\bar{V}~{}i\bar{W}\right]^{T}=0$ with the matrix [99]
@@ A.2 Solutions in piezoelectric materials
where the electrical field $E_{k}$ is applied along the x-axis. Assume the electrical potential is $\Phi$, then the electrical field satisfies that $E_{k}=-\partial\Phi/\partial x_{k}$. Then the motion equation within the piezoelectric materials reads
@@
Then the motion of the piezoelectric material reads
@@
$\rho\frac{\partial^{2}u_{x}}{\partial t^{2}}=$ $c_{11}\frac{\partial^{2}u_{x}}{\partial x^{2}}+c_{44}\left[\frac{\partial^{2}u_{x}}{\partial y^{2}}+\frac{\partial^{2}u_{x}}{\partial z^{2}}\right]$ (65)
$+\left(c_{12}+c_{44}\right)\left[\frac{\partial^{2}u_{y}}{\partial x\partial y}+\frac{\partial^{2}u_{z}}{\partial x\partial z}\right]+2e_{14}\frac{\partial^{2}\phi}{\partial y\partial z},$
$\rho\frac{\partial^{2}u_{y}}{\partial t^{2}}=$ $c_{11}\frac{\partial^{2}u_{y}}{\partial y^{2}}+c_{44}\left[\frac{\partial^{2}u_{y}}{\partial x^{2}}+\frac{\partial^{2}u_{y}}{\partial z^{2}}\right]$ (66)
$+\left(c_{12}+c_{44}\right)\left[\frac{\partial^{2}u_{x}}{\partial y\partial x}+\frac{\partial^{2}u_{z}}{\partial y\partial z}\right]+2e_{14}\frac{\partial^{2}\phi}{\partial x\partial z},$
$\rho\frac{\partial^{2}u_{z}}{\partial t^{2}}=$ $c_{11}\frac{\partial^{2}u_{z}}{\partial z^{2}}+c_{44}\left[\frac{\partial^{2}u_{z}}{\partial x^{2}}+\frac{\partial^{2}u_{z}}{\partial y^{2}}\right]$ (67)
$+\left(c_{12}+c_{44}\right)\left[\frac{\partial^{2}u_{x}}{\partial z\partial x}+\frac{\partial^{2}u_{y}}{\partial z\partial y}\right]+2e_{14}\frac{\partial^{2}\phi}{\partial x\partial y},$
$\varepsilon\Delta\phi=2e_{14}\left[\frac{\partial^{2}u_{x}}{\partial y\partial z}+\frac{\partial^{2}u_{y}}{\partial x\partial z}+\frac{\partial^{2}u_{z}}{\partial x\partial y}\right],$ (68)
@@
When the sample is fabricated on the polished (100) surface, and the acoustic wave is propagated along the [011] direction, then the acoustic wave can be evaluated as [46]
$\left\{\begin{array}[]{l}u_{x}=\frac{u}{\sqrt{2}}e^{-kz+ik(x+y)/\sqrt{2}-ivkt},\\ u_{y}=\frac{u}{\sqrt{2}}e^{-kz+ik(x+y)/\sqrt{2}-ivkt},\\ u_{z}=\tilde{u}_{z}e^{-kz+ik(x+y)/\sqrt{2}-ivkt}.\\ \end{array}\right.$ (74)
Then Eq. (68) reads
$\varepsilon\Delta\phi=\begin{cases}2e_{14}\left[2\frac{\partial^{2}u_{x}}{\partial x\partial z}+\frac{\partial^{2}u_{z}}{\partial x^{2}}\right],&0<z<H,\\ 0,&\mbox{else},\end{cases}$ (75)
@@
Then the amplitude of the electric field by SAW can be calculated as [100]
@@ A.3 Electrical fields induced by the surface acoustic wave
$=-\frac{\mu_{\rm eff}}{\sqrt{\left(\frac{E_{\Lambda}}{2}\right)^{2}+\left(\mu_{eff}|\vec{E}|\right)^{2}}}\left[\left(\frac{\partial^{2}\phi}{\partial x^{2}}\right)^{2}+\left(\frac{\partial^{2}\phi}{\partial z^{2}}\right)^{2}+2\left(\frac{\partial^{2}\phi}{\partial z\partial x}\right)^{2}\right]<0,$
where $\phi=\phi(x,z,t)$ is given in Eq. (80) and the inequality holds when $\mu_{\rm eff}>0$ [60].

### v4

@@ IV.2 Cooperative shielding
We first consider a simplified case that $\gamma_{nj}\equiv\gamma$, and then consider a general case. As shown in Fig. 4(b), where $c_{n}(T)$ is for $\gamma\neq 0$ and $c_{n}^{\prime}(T)$ is for $\gamma=0$ [66], the effect of $\gamma$ can be evaluated with $\max_{n}\left|\left|c_{n}(T)\right|^{2}-\left|c_{n}^{\prime}(T)\right|^{2}\right|$. The comparisons in Figs. 4(c) and (d) reveal that the long range interactions among trapped polar molecules can be cancelled out when the number of trapped sites is large based on the designed initial condition.
@@
where $c_{n}(t)$, as a generalization of $p_{n}$ in Eq. (15), represents the amplitude that the $n$th site is occupied. For the single-layer trapped polar molecules with a single SAW layer and externa electrical fields, $N$ can be finite. When the polar molecules are trapped with two SAW layers as in Fig. 3(a), $N$ can be infinitely large.
Given the Hamiltonian as Eq. (23), the evolution of the amplitudes can be written according to the Schrödinger equation $|\dot{\psi}(t)\rangle=-iH_{N}|\psi(t)\rangle$ as [66]
@@
with $\hbar=1$. In Fig. 5, we compare the dynamics of Eq. (25) with different number of sites and different initial conditions. The simulations clarify that long-range interactions can affect the dynamics of the lattices, but it is related to the initial condition and number of sites. If the lattices are initially random or uniformly occupied, the influence of long-range interactions is much smaller or better shielded than the case that a small subset of lattices are initially occupied.
Figure 5: $\left|\left|c_{n}(T)\right|^{2}-\left|c_{n}^{\prime}(T)\right|^{2}\right|$ varies with the number of cites $N$, where $c_{n}(T)$ represents the terminal values in Eq. (25) with $\gamma_{nj}=0$ at $T=10$s, and $c_{n}^{\prime}(T)$ represents that when $\gamma_{nj}\neq 0$. In all the simulations, $V=1$ and $U_{0}=0.1$. In (a-c), $\gamma_{nj}\equiv\gamma=2$, and in (d-f), $\gamma_{nj}=V|n-j|^{-3}$. In (a) and (d), $c_{n}(0)=1/\sqrt{N}$. In (b) and (e), $c_{n}(t)$ and $c_{n}^{\prime}(t)$ are randomly initialized. In (c) and (f), $c_{1}(0)=c_{2}(0)=1/\sqrt{2}$ and $c_{n}(0)=0$ for $n>2$.
@@ V Bose-Hubbard model for trapped molecules
$\approx\Delta\left(\frac{1-\Delta^{2}}{2m_{0}}+B_{0}e^{-kz}\cos\left[k(x-vt)\right]\right)e^{-\Delta^{2}}$
$\approx\Delta\left(\tilde{B}_{0}e^{-kz}+\delta_{J}\right)e^{-\Delta^{2}}$
@@
where $\delta_{\epsilon}$ is a bounded perturbation similarly as in Eq. (30) within a short time scale. In the following, we first study a simplified case with $\delta_{J}=\delta_{\epsilon}=0$ and then generalize to the case that $\delta_{J},\delta_{\epsilon}\neq 0$.
@@ V.1 Influence by $n_{0}$, uncertainties and thermal effects
Figure 7: (a) Phase transition boundaries when $n_{0}=1,4,7,10,13,16$ and $f_{n_{0}}=0.085786,0.027864,\cdots,0.0075775$, respectively. (b) The parameters of the Bose-Hubbard model affected by random components $\delta_{J}\in[-5,5]$ and $\delta_{\epsilon}\in[-100,100]$. (c) The phase transition between superfluid (red) and Mott insulator (blue) evaluated by $J/U$ with the uncertain parameters in (b). (d) $J/U$ affected by the wave number $k$ when $N=5$ and $\delta_{J}\in[-5,5]$. In all the simulations, $\tilde{B}_{0}=100$, the distance between two nearest neighbor trapped polar molecules is $0.005/N$m, and $k=50$ in (a)-(c).
@@ V.2 Influence by the wave number
Figure 9: The influence on the phase transition by the wave number $k$ of the acoustic wave. (a)-(c) represent the phase transition between the superfluid (red region) and Mott insulator (blue region) with $n_{0}=1$ and the uncertainties in Fig. 7 (b). (d)-(f) represent the phase transition boundaries with $n_{0}=1,4,7,10,13,16$, and different colors represent the regions for the superfluid state separated by the boundaries with different $n_{0}$ compared with the white insulator region. In all the simulations, $B_{0}=100$, the distance between two nearest neighbor trapped polar molecules is $0.005/N$m.
@@ VI Conclusion
We have proposed a method to trap polar molecules with the electrical field induced by surface acoustic waves. Assisted by external electrical fields, the polar molecules can be trapped and arranged into single or multi-parallel layers away from the piezoelectric material. Depending on the uniformity of the layers in the longitudinal direction, the attractive interactions among the molecules in each layer can be different, which can further affect the binding energy of the trapped polar molecule array. For a single layer of trapped polar molecules, the final steady distribution of the molecules can be affected by the trapping approach with finite or infinite lattice sites. The trapped polar molecules can be used to construct the lattice based Bose-Hubbard model. The phase transition between the superfluid and Mott insulator can be controlled by designing the values and spatial distributions of the external electrical force. The advantages of SAW for trapping polar molecules are that its induced electrical field satisfies the condition of stability for trapping a polar particle, the field can be real-time modulated by the electrical field applied to the input IDT, and arbitrary three-dimensional polar molecule lattices can in principle be trapped by tuning another applied electrical force. Our proposal can be more efficient compared with traditional trapping methods with electrical force such as in Refs. [21, 22, 23, 24]. Besides, the exponential decrease of the electrical field produced by SAW along the vertical direction in the open space can induce the transition between Mott insulator and superfluid, and the overall architecture can be fabricated on-chip with extensive possible expansions. From the perspective of experimental realization, the deceleration of the molecule gas has been realized in Refs. [20, 52, 89], the on-chip design of the surface acoustic wave resonator has been demonstrated in Ref. [53], and the method of shaping electrical fields in the open space between two SAW layers has been introduced in Refs. [21, 54]. These technologies support the viability of the proposed trapping method. Our proposal has potential applications in the construction of superfluid and Mott insulator [90, 91], phase transitions [92, 93, 94], lattice QED [95, 74], as well as quantum information processing [96]. This might open an easy way to hybridize molecules with solid state quantum devices, in contrast to the trapping molecules using optical lattices.
Acknowledgement.— Y.X.L. is supported by the National Science Foundation of China (NSFC) under Grants No. 12374483 and No. 92365209. R.B.W. is supported by the NSFC under Grants No. 61833010 and No. 62173201.
@@ A.1 Solutions in isotropic materials
The vector format of the displacement $\mathbf{u}=\left[u_{1},u_{2},u_{3}\right]^{T}$ can be represented as [97]
@@
where $\bar{U}$, $\bar{V}$ and $\bar{W}$ are constants determined by $q$ representing the possible decay in the direction of $x_{3}$, $k$ is the wave number, $v$ is the velocity of the acoustic wave, $l=\cos\theta$, $m=\sin\theta$, $\theta$ is the angle between the propagating direction of the wave front with the $x_{3}$-axis. Thus $l$ and $m$ can represent the propagating direction of the acoustic wave. Take Eq. (45) into Eqs. (42,43,44), we have $\mathbf{P}(q)\left[\bar{U}~{}\bar{V}~{}i\bar{W}\right]^{T}=0$ with the matrix [98]
@@ A.2 Solutions in piezoelectric materials
where the electrical field $E_{k}$ is applied along the x-axis. Assume the electrical potential is $\Phi$, the electrical field satisfies that $E_{k}=-\partial\Phi/\partial x_{k}$. Thus the motion equation within the piezoelectric materials reads
@@
The motion of the piezoelectric material reads
@@
$\rho\frac{\partial^{2}u_{x}}{\partial t^{2}}=$ $c_{11}\frac{\partial^{2}u_{x}}{\partial x^{2}}+c_{44}\left(\frac{\partial^{2}u_{x}}{\partial y^{2}}+\frac{\partial^{2}u_{x}}{\partial z^{2}}\right)$ (65)
$+\left(c_{12}+c_{44}\right)\left(\frac{\partial^{2}u_{y}}{\partial x\partial y}+\frac{\partial^{2}u_{z}}{\partial x\partial z}\right)+2e_{14}\frac{\partial^{2}\phi}{\partial y\partial z},$
$\rho\frac{\partial^{2}u_{y}}{\partial t^{2}}=$ $c_{11}\frac{\partial^{2}u_{y}}{\partial y^{2}}+c_{44}\left(\frac{\partial^{2}u_{y}}{\partial x^{2}}+\frac{\partial^{2}u_{y}}{\partial z^{2}}\right)$ (66)
$+\left(c_{12}+c_{44}\right)\left(\frac{\partial^{2}u_{x}}{\partial y\partial x}+\frac{\partial^{2}u_{z}}{\partial y\partial z}\right)+2e_{14}\frac{\partial^{2}\phi}{\partial x\partial z},$
$\rho\frac{\partial^{2}u_{z}}{\partial t^{2}}=$ $c_{11}\frac{\partial^{2}u_{z}}{\partial z^{2}}+c_{44}\left(\frac{\partial^{2}u_{z}}{\partial x^{2}}+\frac{\partial^{2}u_{z}}{\partial y^{2}}\right)$ (67)
$+\left(c_{12}+c_{44}\right)\left(\frac{\partial^{2}u_{x}}{\partial z\partial x}+\frac{\partial^{2}u_{y}}{\partial z\partial y}\right)+2e_{14}\frac{\partial^{2}\phi}{\partial x\partial y},$
$\varepsilon\Delta\phi=2e_{14}\left(\frac{\partial^{2}u_{x}}{\partial y\partial z}+\frac{\partial^{2}u_{y}}{\partial x\partial z}+\frac{\partial^{2}u_{z}}{\partial x\partial y}\right),$ (68)
@@
When the sample is fabricated on the polished (100) surface, the acoustic wave propagates along the [011] direction, and is of the format [51, 45, 98]
$\left\{\begin{array}[]{l}u_{x}=\frac{u}{\sqrt{2}}e^{-kz+ik(x+y)/\sqrt{2}-ivkt},\\ u_{y}=\frac{u}{\sqrt{2}}e^{-kz+ik(x+y)/\sqrt{2}-ivkt},\\ u_{z}=\tilde{u}_{z}e^{-kz+ik(x+y)/\sqrt{2}-ivkt}.\end{array}\right.$ (74)
Equation (68) reads
$\varepsilon\Delta\phi=\begin{cases}2e_{14}\left(2\frac{\partial^{2}u_{x}}{\partial x\partial z}+\frac{\partial^{2}u_{z}}{\partial x^{2}}\right),&0<z<H,\\ 0,&\mbox{else},\end{cases}$ (75)
@@
The amplitude of the electric field by SAW can be calculated as [99]
@@ A.3 Electrical fields induced by the surface acoustic wave
$=-\frac{\mu_{\rm eff}^{2}}{\sqrt{\left(\frac{E_{\Lambda}}{2}\right)^{2}+\left(\mu_{\rm eff}|\vec{E}|\right)^{2}}}\left[\left(\frac{\partial^{2}\phi}{\partial x^{2}}\right)^{2}+\left(\frac{\partial^{2}\phi}{\partial z^{2}}\right)^{2}+2\left(\frac{\partial^{2}\phi}{\partial z\partial x}\right)^{2}\right]<0,$
where $\phi=\phi(x,z,t)$ is given in Eq. (80[59, 102].